# Susceptibilities of rotating quark matter in Fourier-Bessel basis

### v1

@@ II.3 Boundary conditions in finite cylinder
$R\int_{-\infty}^{\infty}dz\int_{0}^{2\pi}d\theta\,\bar{\psi}\gamma^{r}\psi\Bigl{|}_{r=R}=0.$ (20)
A sufficient condition for Eq. (20) is that the radial momentum $p_{\perp}$ is discretized, as follows [59, 58]:
$p_{\perp}\to p_{l,k}=\begin{cases}\xi_{l,k}/R\quad&{\rm for}\quad l=0,1,\cdots,\\ \xi_{-l-1,k}/R\quad&{\rm for}\quad l=-1,-2,\cdots,\end{cases}$ (21)
@@
$\sum_{l=-\infty}^{\infty}\int\frac{d^{3}p}{(2\pi)^{3}}\to\int\frac{dp_{z}}{2\pi}\sum_{k=1}^{\infty}\sum_{l=-\infty}^{\infty}\frac{1}{\pi R^{2}N_{l,k}^{2}},\qquad N_{l,k}=\begin{cases}|J_{l+1}(\xi_{l,k})|\quad&\text{for}\quad l=0,1,\cdots,\\ |J_{-l}(\xi_{-l-1,k})|\quad&\text{for}\quad l=-1,-2,\cdots\end{cases}$ (22)
We note that the discretized momentum (21) and the resulting normalization factor $N_{l,k}$ are invariant under the replacement $j=l+1/2\leftrightarrow-j=-l-1/2$, or equivalently $l\leftrightarrow-l-1$. This symmetry implies that our boundary condition does not violates time reversal. While in the vacuum the susceptibility functions actually become independent of $\Omega$, the $r$-dependence persists even in the vacuum due to the requirement to satisfy the boundary condition of the no-incoming flux condition in Eq. (20). This behavior will be demonstrated in detail later using a model approach.
@@
$\Delta(x,x^{\prime})=T\sum_{n=-\infty}^{\infty}\int\frac{dp_{z}}{2\pi}\sum_{k=1}^{\infty}\sum_{l=-\infty}^{\infty}e^{-i\omega_{n}(\tau-\tau^{\prime})+ip_{z}(z-z^{\prime})}\frac{\mathcal{S}_{l}(-i\omega_{n},p_{l,k},p_{z};\bm{x}_{\perp},\bm{x}_{\perp}^{\prime})}{\pi R^{2}N_{l,k}^{2}},$ (23)
[…]
III Susceptibility functions within an NJL model
@@
${\cal L}_{\rm int}=\frac{g_{s}+g_{d}}{2}\left[(\bar{q}q)^{2}+\sum_{a=1}^{3}(\bar{q}i\gamma_{5}\tau^{a}q)^{2}\right]+\frac{g_{s}-g_{d}}{2}\left[\sum_{a=1}^{3}(\bar{q}\tau^{a}q)^{2}+(\bar{q}i\gamma_{5}q)^{2}\right],$ (25)
where $\tau^{i=1,2,3}$ are Pauli matrices with $\tau^{0}={\bm{1}}_{2\times 2}$, and $g_{s}$ and $g_{d}$ are the coupling constants. The four-point interaction terms characterized by $g_{s}$ are invariant parts under the global $U(2)_{L}\times U(2)_{R}$ chiral symmetry. On the other hand, the interaction described by $g_{d}$ breaks the $U(1)_{A}$ symmetry and corresponds to ‘t Hooft-Kobayashi-Maskawa determinant term [60, 61, 62, 63].
Before performing explicit calculations of the susceptibility function, we review the Ward-Takahashi (WT) identity associated with the chiral symmetry [64, 48, 49]. This identity serves as a theoretical constraint and a guiding principle for the evaluation of the susceptibility function within the present model analysis.
@@
$\langle\mathcal{O}(y)\rangle=\frac{1}{Z}\int\mathcal{D}\bar{q}\mathcal{D}q\,\mathcal{O}(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q},q)\right],\quad Z=\int\mathcal{D}\bar{q}\mathcal{D}q\,\exp\left[\int d^{4}x^{\prime}\,\mathcal{L}(\bar{q},q)\right].$ (26)
@@
$q^{\prime}=(1+i\epsilon^{i}\gamma_{5}\tau^{i}/2)q,\quad\bar{q}^{\prime}=\bar{q}(1+i\epsilon^{i}\gamma_{5}\tau^{i}/2),$ (27)
@@
$\int\mathcal{D}\bar{q}^{\prime}\mathcal{D}q^{\prime}\,\bar{q}^{\prime}(y)i\gamma_{5}\tau^{i}q^{\prime}(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q}^{\prime},q^{\prime})\right]=\int\mathcal{D}\bar{q}\mathcal{D}q\,\bar{q}(y)i\gamma_{5}\tau^{i}q(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q},q)\right]$ (28)
Under the transformation (27), no nontrivial Jacobian emerges in the functional measure, but the Lagrangian and the pseudoscalar change as
$\mathcal{L}(\bar{q}^{\prime},q^{\prime})=\mathcal{L}(\bar{q},q)-m_{0}\epsilon^{i}\bar{q}(x)i\gamma_{5}\tau^{i}q(x),\quad\bar{q}^{\prime}(x)i\gamma_{5}\tau^{i}q^{\prime}(x)=\bar{q}(x)i\gamma_{5}\tau^{i}q(x)-\epsilon^{i}\bar{q}(x)q(x).$ (29)
Plugging them into Eq. (28) and comparing the coefficients of $\epsilon^{i}$ in both sides, we can obtain the desired WT identity,
$\langle\bar{q}q\rangle(x)=-m_{0}\chi_{\pi}(x),$ (30)
@@
$\chi_{\pi}(x)=\int d^{4}x^{\prime}\langle T_{\tau}\bar{q}(x)i\gamma_{5}\tau^{i}q(x)\bar{q}(x^{\prime})i\gamma_{5}\tau^{i}q(x^{\prime})\rangle.$ (31)
Note that the isospin index $i$ is not summed over and the results for $i=1,2,3$ are identical due to the isospin symmetry of the NJL model. Through the WT identity in Eq. (30) the pion susceptibility is directly related to the quark condensate. A crucial remark here is that the above argument is irrelevant to coordinates and boundary conditions, as we have not specified it above. Hence, the WT identity (30) should be respected even in our rotating system.
@@
${\cal L}_{\rm int}\to(g_{s}+g_{d})\langle\bar{q}q\rangle\bar{q}q-\frac{g_{s}+g_{d}}{2}\langle\bar{q}q\rangle^{2}+{\cal L}_{\rm int}.$ (32)
Here, the interaction term ${\cal L}_{\rm int}$ is interpreted as a fluctuation around the vacuum in which spontaneous chiral symmetry breaking occurs. Under the mean-field approximation, the effective action and the partition function are described by
@@
[…]
While in general this cannot be solved without specifying the inhomogeneity of the mass, under the local density approximation the solution is represented as the same form as the free theory, up to $O(\partial_{r}M/M^{2})$. That is, the propagator is given by Eq. (23) but with the replacement of the mass, as follows:
$\Delta_{\mathrm{q}}(x,x^{\prime};M(r))=\Delta(x,x^{\prime})\Big{|}_{m\to M(r)}\,,$ (35)
@@
$\Delta_{\mathrm{q}}(x^{\prime},x;M(r^{\prime}))=-\gamma^{0}\Delta_{\mathrm{q}}^{\dagger}(x,x^{\prime};M(r))\gamma^{0}.$ (36)
Plugging the solution (35) into this relation, we can arrive at
$\Delta_{\mathrm{q}}(x^{\prime},x;M(r^{\prime}))=\Delta_{\mathrm{q}}(x^{\prime},x;M(r)).$ (37)
This relation is advantageous in evaluating two-point correlations of quark-bilinears, which we analyze in the following sections. Thanks to Eq. (37), such correlators involving the product of two propagators are described with only one mass, enabling us to employ the same computational manner as that of the free theory.
III.1 Meson susceptibilities
In this subsection, we present the susceptibility functions using the quark (Matsubara) propagator based on the Fourier-Bessel basis, under both the mean-field approximation and the local density approximation. In a similar manner to that in the subsection II.2, we introduce the spurion term
$\begin{split}{\cal L}_{\rm sp}&=\zeta_{\sigma}{\cal O}_{\sigma}+\zeta_{\pi}{\cal O}_{\pi}+\zeta_{\delta}{\cal O}_{\delta}+\zeta_{\eta}{\cal O}_{\eta},\end{split}$ (38)
@@ III.1 Meson susceptibilities
${\cal O}_{\sigma}=\bar{q}q,\quad{\cal O}_{\pi}=\bar{q}i\gamma_{5}\tau^{3}q,\quad{\cal O}_{\eta}=\bar{q}i\gamma_{5}q,\quad{\cal O}_{\delta}=\bar{q}\tau^{3}q.$ (39)
Note that we solely focus on the third isospin component of the quark bilinears corresponding to the pion and delta, as isospin symmetry guarantees equivalent results for the first and second components. By including it in the NJL Lagrangian in Eq. (24), the susceptibility functions can be defined as
$\begin{split}\int d^{4}x\chi_{\alpha}(x)&=\frac{\partial^{2}\Gamma}{\partial\zeta_{\alpha}^{2}}\Biggl{|}_{\zeta_{\alpha}=0}=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle,\qquad\alpha=\pi,\eta,\sigma,\delta.\end{split}$ (40)
@@
$\begin{split}&\int d^{4}x\,\chi_{\pi}\left(r;M(r)\right)=\int d^{4}x\int d^{4}x^{\prime}\left\langle T_{\tau}{\cal O}_{\pi}(x)\,{\cal O}_{\pi}(x^{\prime})\exp\left[\int d^{4}x_{1}\frac{G_{\pi}}{2}(\bar{q}i\gamma_{5}\tau^{k}q)^{2}\right]\right\rangle_{0},\end{split}$ (41)
with $G_{\pi}=g_{s}+g_{d}$ and $\langle\cdots\rangle_{0}$ representing the vacuum expectation value under the mean-field approximation without the fluctuating interaction term ${\cal L}_{\rm int}$. Furthermore, we have extracted the relevant four-point quark interaction term for the pion susceptibility, characterized by $G_{\pi}$. To proceed with the evaluation of $\chi_{\pi}$, we will perform an expansion in terms of $G_{\pi}$. The leading term of the $G_{\pi}$ expansion in Eq. (41) is identified as the one-loop result of the pion susceptibility $\chi_{\pi}^{\rm(1-loop)}\left(r;M(r)\right)$, which is written as follows:
$\begin{split}\int d^{4}x\chi_{\pi}^{\rm(1-loop)}\left(r;M(r)\right)&=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\pi}(x)\,{\cal O}_{\pi}(x^{\prime})\rangle_{0}\\ &=-N_{c}N_{f}\int d^{4}x\int d^{4}x^{\prime}{\rm tr}\left[\Delta_{\mathrm{q}}\left(x,x^{\prime};M(r)\right)i\gamma_{5}\Delta_{\mathrm{q}}\left(x^{\prime},x;M(r^{\prime})\right)i\gamma_{5}\right],\end{split}$ (42)
@@
To proceed with the evaluation, we mention two important remarks. The first one is about the relation between the integrands in both sides. While in translational invariant systems, this equation would provides the equality between two integrands, in the present case the difference between them are, without loss of generality, described by an $r$-dependent redundant $\alpha_{1}(r)$, whose radial integral vanishes. Another is that thank to the relation (37), the integration over $x^{\prime}$ as the usual fermionic one-loop diagram with the mass $M(r)$. Therefore, we obtain
$\chi_{\pi}^{({\rm 1-loop})}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{1}{E}\left(f(p_{z};\Lambda)-2n_{F}(E)\right)+\alpha_{1}(r)$ (43)
@@
$f(p_{z};\Lambda)=\frac{\Lambda^{10}}{\Lambda^{10}+\left(\sqrt{p_{l,k}^{2}+p_{z}^{2}}\right)^{10}},$ (44)
with $\Lambda$ being the UV cutoff. As for the rotational effect, its explicit dependence appears only in the Fermi distribution function. At zero temperature, however, the distribution function vanishes due to the infrared energy gap and the causality constraint, at least for the present discretized momentum (21). In other words, there is no visible rotational effect on thermodynamic quantities, as shown in Ref. [7]. The same feature is also found for a different boundary condition [8] and for a rotating sphere [65, 21].
@@
$\begin{split}\int d^{4}x\chi_{\pi}(r;M(r))&=\int d^{4}x\left[\left\{\chi_{\pi}^{\rm(1-loop)}+G_{\pi}\left(\chi_{\pi}^{\rm(1-loop)}\right)^{2}+G_{\pi}^{2}\left(\chi_{\pi}^{\rm(1-loop)}\right)^{3}+\cdots\right\}+\mathcal{R}\right]\\ &=\int d^{4}x\frac{\chi_{\pi}^{({\rm 1-loop})}}{1-G_{\pi}\chi_{\pi}^{({\rm 1-loop})}}+\int d^{4}x\mathcal{R},\end{split}$ (45)
where $\mathcal{R}=\mathcal{R}(\alpha_{1}(r),\alpha_{2}(r),\cdots)$ is the total redundant contribution determined by $\alpha_{n}(r)$ being the parameter emerging in the $O(G_{\pi}^{n})$ computation. The pion susceptibility is now represented in the bubble ring diagram form, as in the vacuum case [66], accompanied by $\mathcal{R}$. Indeed, when attempting to satisfy the WT identity in Eq. (30) under both the mean-field approximation and the local density approximation, the redundant term must vanish. That is, by respecting the WT identity, the pion susceptibility is given by the following expression:
$\begin{split}\chi_{\pi}\left(r;M(r)\right)=\frac{\chi_{\pi}^{({\rm 1-loop})}}{1-G_{\pi}\chi_{\pi}^{({\rm 1-loop})}}.\end{split}$ (46)
A detailed verification of whether this pion susceptibility satisfies the WT identity within the current NJL model analysis will be examined at the beginning of section IV.
Besides, using the expression of the pion susceptibility in Eq. (46) as a basis, we proceed to derive the explicit forms of the susceptibility functions for other meson channels. Through the $SU(2)_{A}$ and $U(1)_{A}$ transformations, the quark bilinears in Eq. (39) are transformed into different quark bilinears as 111 These transformations are performed in the context of bilinear field operators $\mathcal{O}_{\alpha}$, but the same results are also derived in the path integral formalism. That is, WT identities similar to Eq. (30) are obtained through the variable transformation, and these identities requires Eq. (48).
$\begin{split}{\cal O}_{\pi}\overset{SU(2)_{A}}{\rightarrow}{\cal O}_{\sigma}\overset{U(1)_{A}}{\rightarrow}{\cal O}_{\eta}\overset{SU(2)_{A}}{\rightarrow}{\cal O}_{\delta}.\end{split}$ (47)
By taking into account the transformations with the result of the pion susceptibility in Eq. (46), the meson susceptibilities, including those of other mesonic channels, are described as follows:
$\begin{split}\chi_{\alpha}\left(r;M(r)\right)&=\frac{\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle_{0}}{1-G_{\alpha}\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle_{0}}\\ &=\frac{\chi_{\alpha}^{({\rm 1-loop})}}{1-G_{\alpha}\chi_{\alpha}^{({\rm 1-loop})}},\qquad\alpha=\pi,\eta,\sigma,\delta\end{split}$ (48)

### v2

@@ 2.3 Boundary conditions in finite cylinder
$R\int_{-\infty}^{\infty}dz\int_{0}^{2\pi}d\theta\,\bar{\psi}\gamma^{r}\psi\Bigl|_{r=R}=0.$ (2.20)
A sufficient condition for Eq. (2.20) is that the radial momentum $p_{\perp}$ is discretized, as follows Hortacsu:1980kv; Ambrus:2015lfr:
$p_{\perp}\to p_{l,k}=\begin{cases}\xi_{l,k}/R\quad&{\rm for}\quad l=0,1,\cdots,\\ \xi_{-l-1,k}/R\quad&{\rm for}\quad l=-1,-2,\cdots,\end{cases}$ (2.21)
@@
$\sum_{l=-\infty}^{\infty}\int\frac{d^{3}p}{(2\pi)^{3}}\to\int\frac{dp_{z}}{2\pi}\sum_{k=1}^{\infty}\sum_{l=-\infty}^{\infty}\frac{1}{\pi R^{2}N_{l,k}^{2}},\qquad N_{l,k}=\begin{cases}|J_{l+1}(\xi_{l,k})|,&l=0,1,\cdots,\\ |J_{-l}(\xi_{-l-1,k})|,&l=-1,-2,\cdots.\end{cases}$ (2.22)
We note that the discretized momentum (2.21) and the resulting normalization factor $N_{l,k}$ are invariant under the replacement $j=l+1/2\leftrightarrow-j=-l-1/2$, or equivalently $l\leftrightarrow-l-1$. This symmetry implies that our boundary condition does not violates time reversal. While in the vacuum the susceptibility functions actually become independent of $\Omega$, the $r$-dependence persists even in the vacuum due to the requirement to satisfy the boundary condition of the no-incoming flux condition in Eq. (2.20). This behavior will be demonstrated in detail later using a model approach.
@@
$\Delta(x,x^{\prime})=T\sum_{n=-\infty}^{\infty}\int\frac{dp_{z}}{2\pi}\sum_{k=1}^{\infty}\sum_{l=-\infty}^{\infty}e^{-i\omega_{n}(\tau-\tau^{\prime})+ip_{z}(z-z^{\prime})}\frac{\mathcal{S}_{l}(-i\omega_{n},p_{l,k},p_{z};\bm{x}_{\perp},\bm{x}_{\perp}^{\prime})}{\pi R^{2}N_{l,k}^{2}},$ (2.23)
with $\omega_{n}=(2n+1)\pi T$ and $\beta=1/T$. In the subsequent sections, the imaginary-time formalism is employed with a focus on finite-temperature systems. Throughout this discussion, spacetime integrals are expressed with the following shorthand notation, $\int d^{4}x=\int_{0}^{\beta}d\tau\int d^{3}x$.
One can employ alternative boundary condition, such as the MIT boundary condition Ambrus:2015lfr; Chernodub:2016kxh; Zhang:2020hct. While different boundary conditions generally lead to quantitative differences in the evaluation of physical quantities, the MIT boundary condition shares several common features with those under our boundary condition. First, under the MIT boundary condition, the rotational effect on thermodynamic quantities at zero temperature, similar to ours. Also, at high temperatures, the boundary effect is negligible compared to the thermal wavelength, $1/T\ll R$ Ebihara:2016fwa, whereas at low temperatures the boundary effect becomes significant. We note that the above formulation is essentially unchanged under different boundary conditions, except for the momentum discretization.
3 Susceptibility functions within an NJL model
@@
${\cal L}_{\rm int}=\frac{g_{s}+g_{d}}{2}\left[(\bar{q}q)^{2}+\sum_{a=1}^{3}(\bar{q}i\gamma_{5}\tau^{a}q)^{2}\right]+\frac{g_{s}-g_{d}}{2}\left[\sum_{a=1}^{3}(\bar{q}\tau^{a}q)^{2}+(\bar{q}i\gamma_{5}q)^{2}\right],$ (3.2)
where $\tau^{i=1,2,3}$ are Pauli matrices with $\tau^{0}={\bm{1}}_{2\times 2}$, and $g_{s}$ and $g_{d}$ are the coupling constants. The four-point interaction terms characterized by $g_{s}$ are invariant parts under the global $U(2)_{L}\times U(2)_{R}$ chiral symmetry. On the other hand, the interaction described by $g_{d}$ breaks the $U(1)_{A}$ symmetry and corresponds to ‘t Hooft-Kobayashi-Maskawa determinant term Kobayashi:1970ji; Kobayashi:1971qz; tHooft:1976rip; tHooft:1976snw.
Before performing explicit calculations of the susceptibility function, we review the Ward-Takahashi (WT) identity associated with the chiral symmetry GomezNicola:2013pgq; GomezNicola:2016ssy; GomezNicola:2017bhm. This identity serves as a theoretical constraint and a guiding principle for the evaluation of the susceptibility function within the present model analysis.
@@
$\langle\mathcal{O}(y)\rangle=\frac{1}{Z}\int\mathcal{D}\bar{q}\mathcal{D}q\,\mathcal{O}(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q},q)\right],\quad Z=\int\mathcal{D}\bar{q}\mathcal{D}q\,\exp\left[\int d^{4}x^{\prime}\,\mathcal{L}(\bar{q},q)\right].$ (3.3)
@@
$q^{\prime}=(1+i\epsilon^{i}\gamma_{5}\tau^{i}/2)q,\quad\bar{q}^{\prime}=\bar{q}(1+i\epsilon^{i}\gamma_{5}\tau^{i}/2),$ (3.4)
@@
$\int\mathcal{D}\bar{q}^{\prime}\mathcal{D}q^{\prime}\,\bar{q}^{\prime}(y)i\gamma_{5}\tau^{i}q^{\prime}(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q}^{\prime},q^{\prime})\right]=\int\mathcal{D}\bar{q}\mathcal{D}q\,\bar{q}(y)i\gamma_{5}\tau^{i}q(y)\exp\left[\int d^{4}x\,\mathcal{L}(\bar{q},q)\right]$ (3.5)
Under the transformation (3.4), no nontrivial Jacobian emerges in the functional measure, but the Lagrangian and the pseudoscalar change as
$\mathcal{L}(\bar{q}^{\prime},q^{\prime})=\mathcal{L}(\bar{q},q)-m_{0}\epsilon^{i}\bar{q}(x)i\gamma_{5}\tau^{i}q(x),\quad\bar{q}^{\prime}(x)i\gamma_{5}\tau^{i}q^{\prime}(x)=\bar{q}(x)i\gamma_{5}\tau^{i}q(x)-\epsilon^{i}\bar{q}(x)q(x).$ (3.6)
Plugging them into Eq. (3.5) and comparing the coefficients of $\epsilon^{i}$ in both sides, we can obtain the desired WT identity,
$\langle\bar{q}q\rangle(x)=-m_{0}\chi_{\pi}(x),$ (3.7)
@@
$\chi_{\pi}(x)=\int d^{4}x^{\prime}\langle T_{\tau}\bar{q}(x)i\gamma_{5}\tau^{i}q(x)\bar{q}(x^{\prime})i\gamma_{5}\tau^{i}q(x^{\prime})\rangle.$ (3.8)
Note that the isospin index $i$ is not summed over and the results for $i=1,2,3$ are identical due to the isospin symmetry of the NJL model. Through the WT identity in Eq. (3.7) the pion susceptibility is directly related to the quark condensate. A crucial remark here is that the above argument is irrelevant to coordinates and boundary conditions, as we have not specified it above. Hence, the WT identity (3.7) should be respected even in our rotating system.
@@
${\cal L}_{\rm int}\to(g_{s}+g_{d})\langle\bar{q}q\rangle\bar{q}q-\frac{g_{s}+g_{d}}{2}\langle\bar{q}q\rangle^{2}+{\cal L}_{\rm int}.$ (3.9)
Here, the interaction term ${\cal L}_{\rm int}$ is interpreted as a fluctuation around the vacuum in which spontaneous chiral symmetry breaking occurs. Under the mean-field approximation, the effective action and the partition function are described by $\Gamma=\ln Z$ with
@@
where the three-dimensional spatial integral is given by $\int d^{3}x=\int_{0}^{R}dr\,r\int_{0}^{2\pi}d\theta\int_{-\infty}^{\infty}dz$, and $M$ denotes the dynamical quark mass, $M=m_{0}-(g_{s}+g_{d})\langle\bar{q}q\rangle$. In the rotating cylindrical system, the quark condensate depends on the radial coordinate $r$ as a consequence of the translational noninvariance, and so does the dynamical mass $M=M(r)$. This effective Lagrangian implies that the quark Matsubara propagator $\Delta_{\mathrm{q}}(x,x^{\prime})=\langle 0|T_{\tau}q(x)\bar{q}(x^{\prime})|0\rangle$ with the imaginary time ordering $T_{\tau}$ is the solution of the following equation:
@@
While in general this cannot be solved without specifying the inhomogeneity of the mass, under the local density approximation the solution is represented as the same form as the free theory, up to $O(\partial_{r}M/M^{2})$. That is, the propagator is given by Eq. (2.23) but with the replacement of the mass, as follows:
$\Delta_{\mathrm{q}}(x,x^{\prime};M(r))=\Delta(x,x^{\prime})\Big|_{m\to M(r)}\,$, (3.12)
@@
$\Delta_{\mathrm{q}}(x^{\prime},x;M(r^{\prime}))=-\gamma^{0}\Delta_{\mathrm{q}}^{\dagger}(x,x^{\prime};M(r))\gamma^{0}.$ (3.13)
Plugging the solution (3.12) into this relation, we can arrive at
$\Delta_{\mathrm{q}}(x^{\prime},x;M(r^{\prime}))=\Delta_{\mathrm{q}}(x^{\prime},x;M(r)).$ (3.14)
This relation is advantageous in evaluating two-point correlations of quark-bilinears, which we analyze in the following sections. Thanks to Eq. (3.14), such correlators involving the product of two propagators are described with only one mass, enabling us to employ the same computational manner as that of the free theory.
3.1 Meson susceptibilities
In this subsection, we present the susceptibility functions using the quark (Matsubara) propagator based on the Fourier-Bessel basis, under both the mean-field approximation and the local density approximation. In a similar manner to that in the subsection 2.2, we introduce the spurion term
$\begin{split}{\cal L}_{\rm sp}&=\zeta_{\sigma}{\cal O}_{\sigma}+\zeta_{\pi}{\cal O}_{\pi}+\zeta_{\delta}{\cal O}_{\delta}+\zeta_{\eta}{\cal O}_{\eta},\end{split}$ (3.15)
@@ 3.1 Meson susceptibilities
${\cal O}_{\sigma}=\bar{q}q,\quad{\cal O}_{\pi}=\bar{q}i\gamma_{5}\tau^{3}q,\quad{\cal O}_{\eta}=\bar{q}i\gamma_{5}q,\quad{\cal O}_{\delta}=\bar{q}\tau^{3}q.$ (3.16)
Note that we solely focus on the third isospin component of the quark bilinears corresponding to the pion and delta, as isospin symmetry guarantees equivalent results for the first and second components. By including it in the NJL Lagrangian in Eq. (3.1), the susceptibility functions can be defined as
$\begin{split}\int d^{4}x\chi_{\alpha}(x)&=\frac{\partial^{2}\Gamma}{\partial\zeta_{\alpha}^{2}}\Biggl|_{\zeta_{\alpha}=0}=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle,\qquad\alpha=\pi,\eta,\sigma,\delta.\end{split}$ (3.17)
@@
$\begin{split}&\int d^{4}x\,\chi_{\pi}\left(r;M(r)\right)=\int d^{4}x\int d^{4}x^{\prime}\left\langle T_{\tau}{\cal O}_{\pi}(x)\,{\cal O}_{\pi}(x^{\prime})\exp\left[\int d^{4}x_{1}\frac{G_{\pi}}{2}(\bar{q}i\gamma_{5}\tau^{k}q)^{2}\right]\right\rangle_{0},\end{split}$ (3.18)
with $G_{\pi}=g_{s}+g_{d}$ and $\langle\cdots\rangle_{0}$ representing the vacuum expectation value under the mean-field approximation without the fluctuating interaction term ${\cal L}_{\rm int}$. Furthermore, we have extracted the relevant four-point quark interaction term for the pion susceptibility, characterized by $G_{\pi}$. To proceed with the evaluation of $\chi_{\pi}$, we will perform an expansion in terms of $G_{\pi}$. The leading term of the $G_{\pi}$ expansion in Eq. (3.18) is identified as the one-loop result of the pion susceptibility $\chi_{\pi}^{\rm(1-loop)}\left(r;M(r)\right)$, which is written as follows:
$\begin{split}&\int d^{4}x\chi_{\pi}^{\rm(1-loop)}\left(r;M(r)\right)\\ &=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\pi}(x)\,{\cal O}_{\pi}(x^{\prime})\rangle_{0}\\ &=-N_{c}N_{f}\int d^{4}x\int d^{4}x^{\prime}{\rm tr}\left[\Delta_{\mathrm{q}}\left(x,x^{\prime};M(r)\right)i\gamma_{5}\Delta_{\mathrm{q}}\left(x^{\prime},x;M(r^{\prime})\right)i\gamma_{5}\right],\end{split}$ (3.19)
@@
To proceed with the evaluation, we mention two important remarks. The first one is about the relation between the integrands in both sides. While in translational invariant systems, this equation would provides the equality between two integrands, in the present case the difference between them are, without loss of generality, described by an $r$-dependent redundant $\alpha_{1}(r)$, whose radial integral vanishes. Another is that thank to the relation (3.14), the integration over $x^{\prime}$ as the usual fermionic one-loop diagram with the mass $M(r)$. Therefore, we obtain
$\chi_{\pi}^{({\rm 1-loop})}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{f(p_{z};\Lambda)-2n_{F}(E)}{E}+\alpha_{1}(r)$ (3.20)
@@
$f(p_{z};\Lambda)=\frac{\Lambda^{10}}{\Lambda^{10}+\left(\sqrt{p_{l,k}^{2}+p_{z}^{2}}\right)^{10}},$ (3.21)
with $\Lambda$ being the UV cutoff. As for the rotational effect, its explicit dependence appears only in the Fermi distribution function. At zero temperature, however, the distribution function vanishes due to the infrared energy gap and the causality constraint, at least for the present discretized momentum (2.21). In other words, there is no visible rotational effect on thermodynamic quantities, as shown in Ref. Ebihara:2016fwa. The same feature is also found for a different boundary condition Chernodub:2016kxh and for a rotating sphere Zhang:2020hct; Mameda:2023sst.
@@
$\begin{split}\int d^{4}x\chi_{\pi}(r;M(r))&=\int d^{4}x\left[\left\{\chi_{\pi}^{\rm(1-loop)}+G_{\pi}\left(\chi_{\pi}^{\rm(1-loop)}\right)^{2}+G_{\pi}^{2}\left(\chi_{\pi}^{\rm(1-loop)}\right)^{3}+\cdots\right\}+\mathcal{R}\right]\\ &=\int d^{4}x\frac{\chi_{\pi}^{({\rm 1-loop})}}{1-G_{\pi}\chi_{\pi}^{({\rm 1-loop})}}+\int d^{4}x\mathcal{R},\end{split}$ (3.22)
where $\mathcal{R}=\mathcal{R}(\alpha_{1}(r),\alpha_{2}(r),\cdots)$ is the total redundant contribution determined by $\alpha_{n}(r)$ being the parameter emerging in the $O(G_{\pi}^{n})$ computation. The pion susceptibility is now represented in the bubble ring diagram form, as in the vacuum case Hatsuda:1994pi, accompanied by $\mathcal{R}$. Indeed, when attempting to satisfy the WT identity in Eq. (3.7) under both the mean-field approximation and the local density approximation, the redundant term must vanish. That is, by respecting the WT identity, the pion susceptibility is given by the following expression:
$\begin{split}\chi_{\pi}\left(r;M(r)\right)=\frac{\chi_{\pi}^{({\rm 1-loop})}}{1-G_{\pi}\chi_{\pi}^{({\rm 1-loop})}}.\end{split}$ (3.23)
A detailed verification of whether this pion susceptibility satisfies the WT identity within the current NJL model analysis will be examined at the beginning of section 4.
Besides, using the expression of the pion susceptibility in Eq. (3.23) as a basis, we proceed to derive the explicit forms of the susceptibility functions for other meson channels. Through the $SU(2)_{A}$ and $U(1)_{A}$ transformations, the quark bilinears in Eq. (3.16) are transformed into different quark bilinears as #1#1#1These transformations are performed in the context of bilinear field operators $\mathcal{O}_{\alpha}$, but the same results are also derived in the path integral formalism. That is, WT identities similar to Eq. (3.7) are obtained through the variable transformation, and these identities requires Eq. (3.25).
$\begin{split}{\cal O}_{\pi}\overset{SU(2)_{A}}{\rightarrow}{\cal O}_{\sigma}\overset{U(1)_{A}}{\rightarrow}{\cal O}_{\eta}\overset{SU(2)_{A}}{\rightarrow}{\cal O}_{\delta}.\end{split}$ (3.24)
By taking into account the transformations with the result of the pion susceptibility in Eq. (3.23), the meson susceptibilities, including those of other mesonic channels, are described as follows:
$\begin{split}\chi_{\alpha}\left(r;M(r)\right)&=\frac{\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle_{0}}{1-G_{\alpha}\int d^{4}x^{\prime}\langle T_{\tau}{\cal O}_{\alpha}(x)\,{\cal O}_{\alpha}(x^{\prime})\rangle_{0}}\\ &=\frac{\chi_{\alpha}^{({\rm 1-loop})}}{1-G_{\alpha}\chi_{\alpha}^{({\rm 1-loop})}},\qquad\alpha=\pi,\eta,\sigma,\delta,\end{split}$ (3.25)